A single dose of a dietary supplement contains 0.2 gram of calcium and 0.2 gram of vitamin C. A single dose of a secondary supplement contains 0.1 gram of calcium and 0.4 gram of vitamin C. If a person wants to take 0.6 gram of calcium and 1.2 grams of vitamin C, how many doses of each supplement should she take?
I'm guessing you have to write a system of equations but I'm not sure how it should be. This is my guess:
(a= calcium, c= vitamin C)
0.2a + 0.2c = ?
0.1a + 0.4c = ?
I don't know if that's anywhere near correct but if it is, I have no clue what the equations should equal to. Please help! Thank you.

Answer by Nate(3500)

0.2a + 0.2c = ?
0.1a + 0.4c = ?
(.2a + .2c) + (.1a + .4c) = .3a + .6c
.3a + .6c
you want .6 grams of calcium
.3(a)
.3(2) = .6
a = 2
you want 1.2 grams of vitamin C
.6(c)
.6(2) = 1.2
c = 2
Two doses from each supplement.